# Bringing Diversity from Diffusion Models to Semantic-Guided Face Asset Generation

### v1

@@ 4.2. Results
To facilitate validating the on-the-fly interactivity of the proposed method, a web application using browser rendering technologies is created. Figure 19 demonstrates the attributes editing function enabled by our method after creation. Users are presented with options to specify ”General Facial Structure” attributes such as age, ethnicity, and gender, and ”Detailed Facial Structure” features including the shape of the face, nose, ears, chin, and more. Users can see the results in the viewport and rotate the camera and light to inspect details of the generated result. From the technical side, the application employs a front-end system for rendering and a back-end system for providing services and hosting the web pages. The front-end system uses the Unity game engine (Unity Technologies, 2022) and is built against the WebGL platform. All the generation requests regarding changes to age, ethnicity, and gender are sent back to the back-end for processing and returned to the front-end as asset files once ready. The mesh object, albedo map, specular map, and displacement map are the major art assets dynamically generated by our models and saved on the cloud storage associated with an ID in our database for later reference. Changes to other facial feature settings are handled locally on the browser, and no new assets need to be generated by the server, so no significant delay is incurred at this stage to ensure an interactive editing experience. During local interactive modification, related assets, such as eye color textures and geometry offset presets, are hosted by the back-end server statically and are only loaded on demand. Changes to these interactively modifiable features are recorded in the database to make the result reproducible. The back-end system is implemented using ASP.NET Core (Microsoft, 2024) and is modularized, so the system’s capabilities can be expanded with ease. Please check our supplementary materials for the demo video recordings and details.
@@ 5. Limitations
While we have made considerable progress towards a production-quality system for semantic-guided generation PBR face asset, there are aspects of our system that can be improved upon.
By sourcing our training data from diffusion models, we expanded the diversity of our training data greatly. However, our method only works for texture maps, as there is currently no controllable high-quality large-scale generative model for face geometry. So, while we can refine the geometry with a single-view face reconstruction method, the diversity of the initial geometry is still limited by what we have in the scanned dataset.
Furthermore, the invisible parts on the UV texture projected from the frontal view portraits, generated by the diffusion model, need to be supplemented by the scanned dataset. This requires a well-distributed scan data. If the generated portrait has no similar labels in the scanned data, for instance, a dark skin texture may not be accurately completed if there’s no dark skin in the scanned database. Therefore, we aim to explore a multi-view consistent texture synthesis method that can produce a complete texture without the need for inpainting.
Since pre-trained diffusion models are used to create the training data, our generation model may inherit any potential biases present in models like Stable Diffusion. Although we ensure that each attribute has balanced training data through manually defined attributes and distributions in data preparation, it is worth discussing how our method might still inherit issues if the pre-trained models do not express certain attributes diversely. Additionally, due to the high quality of the generated faces, misuse and privacy management are also significant challenges.
@@ 6. Conclusion
We have introduced a novel 3D face generative model that allows for semantic control and creates high-quality face models, which is unprecedented in 3DMM of human face. The main contributions that realize our model is a specifically designed normalization network and a disentangled generator that are able to utilize not only the high-quality scanned models but also in-the-wild face models that are vast in quantity and semantic information. Experiments have demonstrated that our model can effectively normalize faces from arbitrary lighting conditions, generate novel face and perform attribute manipulations on the generated 3D face in multiple semantic directions. We believe that the progress our system achieves has great potential use in many applications, including VFX production, customized digital avatars and the generation of synthetic training data for other fundamental computer vision research.
@@ Acknowledgements.
- Abrevaya et al. (2018) Victoria Fernández Abrevaya, Stefanie Wuhrer, and Edmond Boyer. 2018. Multilinear Autoencoder for 3D Face Model Learning. In 2018 IEEE Winter Conference on Applications of Computer Vision, WACV 2018, Lake Tahoe, NV, USA, March 12-15, 2018. IEEE Computer Society, 1–9. https://doi.org/10.1109/WACV.2018.00007
@@
- Booth et al. (2016) James Booth, Anastasios Roussos, Stefanos Zafeiriou, Allan Ponniah, and David Dunaway. 2016. A 3D Morphable Model Learnt from 10, 000 Faces. In 2016 IEEE Conference on Computer Vision and Pattern Recognition, CVPR 2016, Las Vegas, NV, USA, June 27-30, 2016. IEEE Computer Society, 5543–5552. https://doi.org/10.1109/CVPR.2016.598
@@
- Cao et al. (2014) Chen Cao, Yanlin Weng, Shun Zhou, Yiying Tong, and Kun Zhou. 2014. FaceWarehouse: A 3D Facial Expression Database for Visual Computing. IEEE Trans. Vis. Comput. Graph. 20, 3 (2014), 413–425. https://doi.org/10.1109/TVCG.2013.249
@@
- Dai et al. (2020) Hang Dai, Nick E. Pears, William A. P. Smith, and Christian Duncan. 2020. Statistical Modeling of Craniofacial Shape and Texture. Int. J. Comput. Vis. 128, 2 (2020), 547–571. https://doi.org/10.1007/s11263-019-01260-7
@@
- Goodfellow et al. (2014a) Ian Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron Courville, and Yoshua Bengio. 2014a. Generative adversarial nets. In Advances in neural information processing systems. 2672–2680.
- Goodfellow et al. (2014b) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2014b. Generative Adversarial Nets. In NeurIPS 2014. 2672–2680.
@@
- LeGendre et al. (2018) Chloe LeGendre, Kalle Bladin, Bipin Kishore, Xinglei Ren, Xueming Yu, and Paul Debevec. 2018. Efficient multispectral facial capture with monochrome cameras. In ACM SIGGRAPH 2018 Posters. 1–2.
@@
- Li et al. (2020a) Ruilong Li, Karl Bladin, Yajie Zhao, Chinmay Chinara, Owen Ingraham, Pengda Xiang, Xinglei Ren, Pratusha Prasad, Bipin Kishore, Jun Xing, and Hao Li. 2020a. Learning Formation of Physically-Based Face Attributes. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 3407–3416. https://doi.org/10.1109/CVPR42600.2020.00347
- Li et al. (2017) Tianye Li, Timo Bolkart, Michael J. Black, Hao Li, and Javier Romero. 2017. Learning a model of facial shape and expression from 4D scans. ACM Trans. Graph. 36, 6 (2017), 194:1–194:17. https://doi.org/10.1145/3130800.3130813
@@
- Liu et al. (2022) Shichen Liu, Yunxuan Cai, Haiwei Chen, Yichao Zhou, and Yajie Zhao. 2022. Rapid Face Asset Acquisition with Recurrent Feature Alignment. ACM Trans. Graph. 41, 6, Article 214 (nov 2022), 17 pages. https://doi.org/10.1145/3550454.3555509
@@
- Ranjan et al. (2018) Anurag Ranjan, Timo Bolkart, Soubhik Sanyal, and Michael J. Black. 2018. Generating 3D Faces Using Convolutional Mesh Autoencoders. In Computer Vision - ECCV 2018 - 15th European Conference, Munich, Germany, September 8-14, 2018, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 11207), Vittorio Ferrari, Martial Hebert, Cristian Sminchisescu, and Yair Weiss (Eds.). Springer, 725–741. https://doi.org/10.1007/978-3-030-01219-9_43
@@
- Shao et al. (2018) Hang Shao, Abhishek Kumar, and P. Thomas Fletcher. 2018. The Riemannian Geometry of Deep Generative Models. In 2018 IEEE Conference on Computer Vision and Pattern Recognition Workshops, CVPR Workshops 2018, Salt Lake City, UT, USA, June 18-22, 2018. IEEE Computer Society, 315–323. https://doi.org/10.1109/CVPRW.2018.00071
- Shen et al. (2020) Yujun Shen, Jinjin Gu, Xiaoou Tang, and Bolei Zhou. 2020. Interpreting the Latent Space of GANs for Semantic Face Editing. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 9240–9249. https://doi.org/10.1109/CVPR42600.2020.00926
@@
- Smith et al. (2020) William A. P. Smith, Alassane Seck, Hannah Dee, Bernard Tiddeman, Joshua B. Tenenbaum, and Bernhard Egger. 2020. A Morphable Face Albedo Model. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 5010–5019. https://doi.org/10.1109/CVPR42600.2020.00506
@@
- Thong et al. (2023) William Thong, Przemyslaw Joniak, and Alice Xiang. 2023. Beyond Skin Tone: A Multidimensional Measure of Apparent Skin Color. In ICCV.
@@
- Zhou et al. (2024) Mingyuan Zhou, Rakib Hyder, Ziwei Xuan, and Guojun Qi. 2024. UltrAvatar: A Realistic Animatable 3D Avatar Diffusion Model with Authenticity Guided Textures. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 1238–1248.

### v2

@@ 4.2.4. Application
We design an on-the-fly interactive web application based on the proposed method to validate its downstream application value. Figure 23 visualizes the interface for a 3D face asset creation and editing tool. The users are given options to specify ”General Facial Structure” attributes such as age, ethnicity, and gender, and ”Detailed Facial Structure” features including the shape of the face, nose, ears, chin, and more. The web application supports free-viewport manipulation and change of lighting for detailed inspection of the generated results. The application employs a front-end system for rendering and a back-end system for providing services and hosting the web pages. The front-end system uses the Unity game engine (Unity Technologies, 2022) and is built against the WebGL platform. All queries on editing age, ethnicity, and gender are processed by the back-end system, which runs our proposed neural networks to generate the corresponding assets. The mesh object, albedo map, specular map, and displacement map are the major art assets dynamically generated by our models and saved on the cloud storage associated with an ID in our database for later reference. Changes to other facial feature settings are handled locally on the browser, and no new assets are generated by the server. This design eliminates significant delay to ensure an interactive editing experience. In a local session, related assets, such as eye color textures and geometry offset presets, are stored on the back-end server and only loaded on demand. Changes to these modifiable features are updated in the database to allow reproducibility. The back-end system is implemented using ASP.NET Core (Microsoft, 2024) and modularized. Please refer to our supplementary materials for video recordings of the web application.
@@ 5. Limitations
While we have made considerable progress towards a production-quality system for semantic-guided generation of PBR face assets, there are aspects of our system that can be improved upon.
By sourcing our training data from diffusion models, we expanded the diversity of our training data greatly. However, our method only works for texture maps, as there is currently no controllable high-quality large-scale generative model for face geometry. While we can refine the geometry with a single-view face reconstruction method, the diversity of the initial geometries is still limited by the scanned dataset.
Our texture normalization matches diffusion-generated UV textures to the scanned-albedo domain. But its accuracy is limited by the small scanned target domain (200 subjects), which can reduce skin-tone fidelity for rare appearances. We therefore estimate a skin color from manually selected flat regions as additional input. With richer scanned-albedo coverage, leveraging general-purpose intrinsic decomposition as additional de-lighting supervision is a potential way to improve normalization further.
Furthermore, the invisible parts of the UV texture projected from the frontal view portraits, generated by the diffusion model, need to be supplemented by the scanned dataset. This requires a well-distributed scan dataset. If the generated portrait has no similar labels in the scanned data, the skin texture may not be accurately completed. Therefore, we aim to explore a multi-view consistent texture synthesis method that can improve our texture completion design.
Since pre-trained diffusion models are used to create the training data, our generation model may inherit any potential biases present in these models. Although we ensure that each attribute has balanced training data through manually defined attributes and distributions in data preparation, it is worth discussing how our method might still inherit issues if the pre-trained models do not express certain attributes diversely. Additionally, due to the high quality of the generated faces, additional care should be taken to prevent misuse and protect privacy.
@@ 6. Conclusion
We have introduced a novel 3D face generative model that allows for semantic control and creates high-quality face models, which is unprecedented in 3DMM of human faces. The main contributions of our model are a specifically designed normalization network and a disentangled generator, which can leverage not only high-quality scanned models but also in-the-wild face models, which are abundant in quantity and semantic information. Experiments have demonstrated that our model can effectively normalize faces under arbitrary lighting conditions, generate novel faces, and perform attribute manipulations on the generated 3D face in multiple semantic directions. We believe that the progress our system achieves has great potential for use in many applications, including VFX production, customized digital avatars, and the generation of synthetic training data for other fundamental computer vision research.
@@ Acknowledgements.
- Abrevaya et al. (2018) Victoria Fernández Abrevaya, Stefanie Wuhrer, and Edmond Boyer. 2018. Multilinear Autoencoder for 3D Face Model Learning. In 2018 IEEE Winter Conference on Applications of Computer Vision, WACV 2018, Lake Tahoe, NV, USA, March 12-15, 2018. IEEE Computer Society, 1–9. doi:10.1109/WACV.2018.00007
- Amos et al. (2016) Brandon Amos, Bartosz Ludwiczuk, and Mahadev Satyanarayanan. 2016. OpenFace: A general-purpose face recognition library with mobile applications. Technical Report. CMU-CS-16-118, CMU School of Computer Science.
@@
- Booth et al. (2016) James Booth, Anastasios Roussos, Stefanos Zafeiriou, Allan Ponniah, and David Dunaway. 2016. A 3D Morphable Model Learnt from 10, 000 Faces. In 2016 IEEE Conference on Computer Vision and Pattern Recognition, CVPR 2016, Las Vegas, NV, USA, June 27-30, 2016. IEEE Computer Society, 5543–5552. doi:10.1109/CVPR.2016.598
@@
- Cao et al. (2014) Chen Cao, Yanlin Weng, Shun Zhou, Yiying Tong, and Kun Zhou. 2014. FaceWarehouse: A 3D Facial Expression Database for Visual Computing. IEEE Trans. Vis. Comput. Graph. 20, 3 (2014), 413–425. doi:10.1109/TVCG.2013.249
@@
- Chen et al. (2024) Xi Chen, Sida Peng, Dongchen Yang, Yuan Liu, Bowen Pan, Chengfei Lv, and Xiaowei Zhou. 2024. Intrinsicanything: Learning diffusion priors for inverse rendering under unknown illumination. In European Conference on Computer Vision. Springer, 450–467.
@@
- Dai et al. (2020) Hang Dai, Nick E. Pears, William A. P. Smith, and Christian Duncan. 2020. Statistical Modeling of Craniofacial Shape and Texture. Int. J. Comput. Vis. 128, 2 (2020), 547–571. doi:10.1007/s11263-019-01260-7
@@
- Fadaeinejad et al. (2025) Amin Fadaeinejad, Abdallah Dib, Luiz Gustavo Hafemann, Emeline Got, Trevor Anderson, Amaury Depierre, Nikolaus F Troje, Marcus A Brubaker, and Marc-André Carbonneau. 2025. Geometry-Aware Texture Generation for 3D Head Modeling with Artist-driven Control. In Proceedings of the Computer Vision and Pattern Recognition Conference. 6196–6206.
@@
- Goodfellow et al. (2014) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2014. Generative Adversarial Nets. In NeurIPS 2014. 2672–2680.
@@
- Kocsis et al. (2024) Peter Kocsis, Vincent Sitzmann, and Matthias Nießner. 2024. Intrinsic image diffusion for indoor single-view material estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 5198–5208.
@@
- Li et al. (2020a) Ruilong Li, Karl Bladin, Yajie Zhao, Chinmay Chinara, Owen Ingraham, Pengda Xiang, Xinglei Ren, Pratusha Prasad, Bipin Kishore, Jun Xing, and Hao Li. 2020a. Learning Formation of Physically-Based Face Attributes. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 3407–3416. doi:10.1109/CVPR42600.2020.00347
- Li et al. (2017) Tianye Li, Timo Bolkart, Michael J. Black, Hao Li, and Javier Romero. 2017. Learning a model of facial shape and expression from 4D scans. ACM Trans. Graph. 36, 6 (2017), 194:1–194:17. doi:10.1145/3130800.3130813
- Li et al. (2025) Zhibing Li, Tong Wu, Jing Tan, Mengchen Zhang, Jiaqi Wang, and Dahua Lin. 2025. IDArb: Intrinsic Decomposition for Arbitrary Number of Input Views and Illuminations. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=uuef1HP6X7
@@
- Liu et al. (2022) Shichen Liu, Yunxuan Cai, Haiwei Chen, Yichao Zhou, and Yajie Zhao. 2022. Rapid Face Asset Acquisition with Recurrent Feature Alignment. ACM Trans. Graph. 41, 6, Article 214 (nov 2022), 17 pages. doi:10.1145/3550454.3555509
- Luo et al. (2024) Jundan Luo, Duygu Ceylan, Jae Shin Yoon, Nanxuan Zhao, Julien Philip, Anna Frühstück, Wenbin Li, Christian Richardt, and Tuanfeng Wang. 2024. Intrinsicdiffusion: Joint intrinsic layers from latent diffusion models. In ACM SIGGRAPH 2024 Conference Papers. 1–11.
@@
- Ranjan et al. (2018) Anurag Ranjan, Timo Bolkart, Soubhik Sanyal, and Michael J. Black. 2018. Generating 3D Faces Using Convolutional Mesh Autoencoders. In Computer Vision - ECCV 2018 - 15th European Conference, Munich, Germany, September 8-14, 2018, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 11207), Vittorio Ferrari, Martial Hebert, Cristian Sminchisescu, and Yair Weiss (Eds.). Springer, 725–741. doi:10.1007/978-3-030-01219-9_43
@@
- Shao et al. (2018) Hang Shao, Abhishek Kumar, and P. Thomas Fletcher. 2018. The Riemannian Geometry of Deep Generative Models. In 2018 IEEE Conference on Computer Vision and Pattern Recognition Workshops, CVPR Workshops 2018, Salt Lake City, UT, USA, June 18-22, 2018. IEEE Computer Society, 315–323. doi:10.1109/CVPRW.2018.00071
- Shen et al. (2020) Yujun Shen, Jinjin Gu, Xiaoou Tang, and Bolei Zhou. 2020. Interpreting the Latent Space of GANs for Semantic Face Editing. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 9240–9249. doi:10.1109/CVPR42600.2020.00926
@@
- Smith et al. (2020) William A. P. Smith, Alassane Seck, Hannah Dee, Bernard Tiddeman, Joshua B. Tenenbaum, and Bernhard Egger. 2020. A Morphable Face Albedo Model. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2020, Seattle, WA, USA, June 13-19, 2020. IEEE, 5010–5019. doi:10.1109/CVPR42600.2020.00506
@@
- Triplegangers (2021) Triplegangers. 2021. Triplegangers Face Models. https://triplegangers.com/. Online; Accessed: 2021-12-05.
@@
- Zhou et al. (2024a) Mingyuan Zhou, Rakib Hyder, Ziwei Xuan, and Guojun Qi. 2024a. UltrAvatar: A Realistic Animatable 3D Avatar Diffusion Model with Authenticity Guided Textures. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 1238–1248.
- Zhou et al. (2024b) Zhenglin Zhou, Fan Ma, Hehe Fan, Zongxin Yang, and Yi Yang. 2024b. HeadStudio: Text to Animatable Head Avatars with 3D Gaussian Splatting. In ECCV.
@@
Appendix A Details of the Comparisons and User Study
Six short textual prompts, each describing a distinct subject identity in terms of age, gender, and ethnicity, were used to generate 3D human models with each method. These prompts served as the primary input to the generation process and were also shown to participants during evaluation. The full list of prompts included: “Elderly Asian Female,” “Young Adult Western Female,” “Teenage African Male,” “Young Western Male,” “Young Asian Female,” and “Middle-Aged African Male.” All other inputs and settings followed the default configurations or recommended usage described in the original implementations of DreamFace (Zhang et al., 2023a) and Describe3D (Wu et al., 2023).
For each prompt, participants viewed the outputs generated by three different methods. The three outputs were displayed in randomized order. The sequence of prompts was also randomized for each participant.
Each output was rated using two questions on a five-point Likert scale (Very Poor to Excellent): (1) “How well does the model match the description?” and (2) “How would you rate the skin detail quality?” In a small number of trials, a third question was included as an attention check, instructing participants to select a specific response (e.g., “Please select ’Poor’.”). Participants who failed one or more of these checks were excluded from the analysis.
Appendix B Hair and Accessories
To support more complete and realistic digital head models, our system includes hair and common accessories such as eyeglasses, hats, and masks. Hair is modeled using Maya XGen and is attached to a predefined scalp mesh. Since all head models share the same topology, the hair can be transferred across different identities by deforming it according to the shared vertex correspondence.
As for accessories, eyeglasses are rigged individually, and hats and masks are deformed using lattice-based transformations, allowing them to fit varying face geometries.
Appendix C Challenging Cases
The examples in Figure 24 demonstrate that directly processing images generated by general-purpose diffusion models to UV space often leads to significant artifacts, such as malformed teeth, unnatural tone, and local geometric distortions. Some are due to imperfect control over the diffusion model, where the final synthesized image does not align pixel-perfectly with the input geometry, causing mismatches in the UV projection. Others are inherent to the generative process. The model produces undesired color tones due to a lack of structural constraints. Although we manually filtered out abnormal textures when constructing 44,000 UV textures, these issues highlight the need for a generative model specifically trained for texture synthesis.
Appendix D Dataset demographic distribution
The Figure 26 illustrates the distribution of demographic information in our high-quality scanned dataset.